# Optimistic Verifiable Training by Controlling Hardware Nondeterminism

### v2

@@ 4 The Nondeterminism Challenge
Although there are user-side controls for forcing deterministic operations within a single GPU architecture (e.g. NVIDIA H100), these controls do not prevent nondeterminism between GPU architectures (e.g. NVIDIA H100 and V100), where trained models can have similar aggregate performance (e.g. accuracy) yet yield very different predictions, as shown in Figure 2 (Crane, 2018a; NVIDIA, 2022). We now discuss the source of nondeterminism between GPU types.
Floating Point Arithmetic: Computers represent real values using integer and floating-point representations, typically using the IEEE 754 standard (Figure 3). There is a tradeoff between the approximation fidelity and the number of bits used to represent the real values. The chosen precision controls the representable numerical range (e.g. 32-bit floating-point values can represent values between $381.17549435e-$ 38 and $383.40282347e+$ 38). Because computers round to representable floating point values, changing the order in which floating point numbers are accumulated can change the resulting sum (Kahan, 1965; Whitehead & Fit-Florea, 2011).
Example: Floating Point Addition Consider three values $a=0.1$, $b=-0.1$ and $c=0.2$ in FP32. The following two orders of summation, $a+c+b$ vs. $a+b+c$, yield different results: $0.200000017881$ vs. $0.200000002980$.
Example Summation Order FP32 FP32 Rounded to FP16
@@
Table 1: Two examples of floating point accumulation error when rounding arithmetic performed higher precision (e.g. FP32) down to lower precision (e.g. FP16). In the first example, error in the FP32 result does not transfer to the rounded FP16 result. In the second example, the errror in the FP32 result does transfer to the rounded result.
Over the course of the many operations during training, this rounding to a representable floating point value can lead to a large difference in the end result between the $\mathrm{trainer}$ and $\mathrm{auditor}$.
Figure 3: We define rounding to $b$ bits as rounding to the nearest 32-bit floating point that has 0s in the last $32-b$ bits of the mantissa, after accounting for the exponent.
@@
Parallel Computation: In a GPU, a single operation (called a kernel) is executed by thousands of threads in parallel. In NVIDIA terminology, GPUs contain a set of streaming multiprocessors (SMs), which run the thread blocks required for the kernel. At the hardware level, the blocks are divided into warps, or groups of threads, that are assigned to the available cores. Applications partition arithmetic workloads (e.g. batch matrix multiplies), differently depending on the GPU architecture to achieve high performance (NVIDIA, 2022), thus changing the order of floating point operations. Different GPUs have a different number and size of compute units. For instance, the number and tile dimensions of tensor core units (specialized matrix multiplication units on the GPU) on NVIDIA Volta and Turing vs. Ampere and Hopper architectures is different (Sun et al., 2022). As a result, a kernel executed on Volta will result in a different value than a kernel executed on Hopper. Other work uses nondeterministic execution patterns, allowing a worker to perform redundant work when it is faster than waiting on sequentially dependent outputs from a slower parallel processor, to improve efficiency (Merrill & Garland, 2016).
@@
Memory Hierarchy and Variable Delays: The time taken for memory access by each thread depends on the physical location of the data, which can create variable delays (Jooybar et al., 2013; Defour & Collange, 2015; Chou et al., 2020). In accessing data, the GPU memory hierarchy consists of large amounts of high bandwidth memory (HBM) and small amounts of fast SRAM memory allocated to each SM, and maintains an L1 and L2 cache to improve access times for reused data. The caches sizes and access times differ across GPU architectures, which affects warp scheduling. For instance, an NVIDIA A100 has 192KB / 40 MB of L1 and L2 cache memory respectively, while the H100 has 256KB / 50MB respectively (NVIDIA, 2023). Further, the physical locations of individual SMs differ in distance from the cache and HBM memory on chip.
Non Determinism To compute primitives such as GEMMs ($D=A\cdot B+C$), the workhorse of machine learning, GPUs split the work of computing the tiles of $D$ across a thread block (NVIDIA, 2023). The results across threads are accumulated. Due to the varied parallelization and data access patterns across architectures, and the fact that floating point addition is not associative, results across GPU architectures are nondeterministic (Crane, 2018a; NVIDIA, 2022).
Any exact and robust verifiable training method would need to control these sources of non-determinism across GPUs.
@@ 5 Method Overview
Figure 4: Divergence between outputs on two different GPU types (in FP64) for a given function and input can result in different rounding choices when rounding to the nearest FP32. We only wish to log rounding decisions for situations like Case A, where the $\mathrm{auditor}$ should copy the $\mathrm{trainer}$’s rounding choice in order to reach the same value. This requires defining a logging region, determined by a threshold $\tau$,
Here we provide our method for the verification game in Section 3 (Teutsch & Reitwießner, 2019). We (1) provide our approach for handling nondeterminism from floating point accumulation, (2) provide the algorithms for the trainer and auditor to verify training, and (3) discuss two strategies for storage cost reduction.
@@ 5.1 Accumulation Errors Start at Higher Precision Bits
Our key idea is that if nondeterminism of training between GPU types occurs due to floating-point operations, then any error will initially be introduced in the lower bits. Suppose that both $\mathrm{trainer}$ and $\mathrm{auditor}$ train at a higher floating-point (e.g. $b_{tr}=64$) precision than the client’s target model precision (e.g. $b_{m}=32$) and then periodically round (e.g. $b_{r}=32$) after intermediate computation steps (e.g. a convolution layer). One might hope that this will “erase” the errors due to nondeterminism, and prevent them from accumulating. Table 1 (Row 1) provides a toy example of how error in higher precision arithmetic is erased when rounding the result to lower precision.
Unfortunately, simply rounding to the nearest FP32 after each computation during training is not sufficient for eliminating nondeterminism. The problem is due to rounding errors that straddle the rounding boundary. Consider Case A in Figure 4, which shows a divergence in the output of a computation using FP64 on two different GPUs. Because the outputs of GPU 1 and 2 are on different sides of the boundary, rounding to the nearest FP32 results in rounding to different values, introducing error. Table 1 (Row 2) provides a toy example of how error in higher precision arithmetic might transfer to the lower precision rounded result.
@@
What if the $\mathrm{trainer}$ records their rounding choice (e.g. up, down, none) for every intermediate computation? The $\mathrm{auditor}$ could then copy the $\mathrm{trainer}$’s choice, and therefore round to the exact same value and successfully control for nondeterminism. However, as we show in Case B in Figure 4, the $\mathrm{auditor}$ should not copy the $\mathrm{trainer}$’s behavior for every output. If a computation output on GPU 1 is too close to the rounded value, then it is possible that GPU 2 is also close in distance but from the opposite direction. In this case, the $\mathrm{auditor}$ should ignore the $\mathrm{trainer}$’s choice. We therefore need to introduce a threshold $\tau$ under which the $\mathrm{trainer}$ does not record their rounding choice.
@@
The success of this method relies on our ability to upper bound the divergence $d_{div}$ between any two different GPUs for any intermediate computation $f$ (i.e. difference in outputs for the same input). Let $\epsilon_{b}$ represent the distance between two FP32 values, after rounding to $b_{r}$ bits of the mantissa (Figure 3) and controlling for the exponent. We need to select $b_{r}$ and $\tau$ such that $d_{div}<\epsilon_{b_{r}}$ and $d_{div}<2\tau$ (Figure 4). Because the set of possible floating point numbers is finite, there exist optimal bounds for $b_{r}$ and $\tau$. In practice, we find that $b_{r}\leq 32$ and $\tau>0.25\cdot\epsilon_{32}$ are sufficient for standard intermediate computations in neural network training (e.g. convolution, layer norm) in FP64. We experiment with different values for $b_{r}$ in Section 6.
We next discuss the specific functions that our method requires, the algorithm details for both the $\mathrm{trainer}$ and $\mathrm{auditor}$, and our approach for reducing storage costs.
Algorithm 1 $\mathsf{log}$
Input: value $x$, rounding amount $b_{r}$, threshold $\tau$, file $F$
1:  $\exp\leftarrow$ get exponent of $x$
2:  if $|x-\mathsf{rnd}_{b_{r}}(x)|>\exp*\tau$ and $x<\mathsf{rnd}_{b_{r}}(x)$ then
3:     $\mathsf{write}(2,F)$ // log rounding up
4:  else if $|x-\mathsf{rnd}_{b_{r}}(x)|>\exp*\tau$ and $x>\mathsf{rnd}_{b_{r}}(x)$ then
5:     $\mathsf{write}(0,F)$ // log rounding down
6:  else
7:     $\mathsf{write}(1,F)$ // log rounding ignore
8:  end if
@@ 5.2 Primitives
Our work assumes both $\mathrm{trainer}$ and $\mathrm{auditor}$ train models using the IEEE-754 standard for floating-points (Figure 3). In addition to requiring $\mathsf{read}$ and $\mathsf{write}$ disk I/O operations, we define the following functions:
@@
    $\mathsf{rnd}_{b_{r}}(x)$: rounds input $x$ to the nearest float up to $b_{r}$ bits of the mantissa, as shown in Figure 3.
@@
    $\mathsf{rev}(x,b_{r},c)$: reverses rounding of input $x$ based on logging direction $c$. If $x<\mathsf{rnd}_{b_{r}}(x)$ and $c=0$, then return $x$ rounded to the nearest float below $x$ with $b_{r}$ precision. If $x>\mathsf{rnd}_{b_{r}}(x)$ and $c=2$, then return $x$ rounded to the nearest float above $x$ with $b_{r}$ precision. Otherwise, do not correct.
@@
    $\mathsf{threshold}(l,b_{r},b_{tr})$: identifies the optimal threshold to log rounding directions (0 or 2) instead of 1, which the $\mathsf{rev}$ function ignores, based on a binary search procedure described in Section 5.5.
@@
    $\mathsf{hash}_{\mathsf{sha256}}(\theta)$: creates a SHA-256 hash of provided model weights $\theta$ via serialization. The model weights should be in the $\mathrm{client}$’s target precision $b_{m}<b_{tr}$.

### v3

@@ 4 The Nondeterminism Challenge
Example Sum Order FP32 Rounded to FP16
@@
Table 1: Two examples of floating point accumulation error when rounding arithmetic performed higher precision (e.g. FP32) down to lower precision (e.g. FP16). In the second example, the error in the FP32 result transfers to the rounded FP16 result.
Although there are user-side controls for forcing deterministic operations within a single GPU architecture , these controls do not prevent nondeterminism between GPU architectures (e.g., NVIDIA H100 and V100), where trained models can have similar aggregate performance (e.g., accuracy) yet yield very different predictions, as shown in Figure 2 Crane (2018a); NVIDIA (2022). There are three main sources of nondeterminism between GPU types:
1. Floating-Point Arithmetic: Computers represent real values using integer and FP representations, typically the IEEE 754 standard (Figure 5). There is a tradeoff between the approximation fidelity and the # of bits used to represent the real values. The chosen precision controls the representable numerical range (e.g., 32-bit FP values can represent values between $381.17549435e-$ 38 and $383.40282347e+$ 38). Because computers round to representable FP values, changing the order in which FP values are accumulated can change the resulting sum Kahan (1965); Whitehead and Fit-Florea (2011). Over the course of the many operations during training, this can lead to a large difference in the end result between the $\mathrm{trainer}$ and $\mathrm{auditor}$.
@@
2. Parallel Computation: In a GPU, a single operation (called a kernel) is executed by thousands of threads in parallel. GPUs contain a set of streaming multiprocessors (SMs), which run the thread blocks required for the kernel. At the hardware level, these blocks are divided into warps that are assigned to the available cores. Because different GPUs have a different number and size of compute units, applications partition arithmetic workloads (e.g., batch matrix multiplies) differently to achieve high performance NVIDIA (2022), thus changing the order of FP operations.
@@
3. Memory Hierarchy and Variable Delays: The time taken for memory access by each thread depends on the physical location of the data, which can create variable delays Jooybar et al. (2013); Defour and Collange (2015); Chou et al. (2020). The GPU memory hierarchy consists of large amounts of high bandwidth memory (HBM) and small amounts of fast SRAM memory, and maintains an L1 and L2 cache to improve access times. The caches sizes and access times differ across GPU architectures (e.g. an NVIDIA A100 has 192KB / 40 MB of L1/L2 cache memory, while the H100 has 256KB / 50MB). This affects warp scheduling, leading to changes in operation ordering resulting in nondeterminism. Finally, to compute primitives such as GEMMs ($D=A\cdot B+C$), the workhorse of machine learning, GPUs split the work of computing the tiles of $D$ across a thread block NVIDIA (2023), resulting in nondeterminism that a robust verification method needs to control.
@@ 5 Method Overview
Figure 3: Divergence between outputs on two different GPUs (in FP64) for a given function and input can result in different rounding choices when rounding to the nearest FP32. We only wish to log rounding decisions for Case A, where the $\mathrm{auditor}$ should copy the $\mathrm{trainer}$’s rounding choice in order to reach the same value. This requires defining a logging region, determined by a threshold $\tau$,
@@ 5.1 Accumulation Errors Start at Higher Precision Bits
Our key idea is that if nondeterminism of training between GPU types occurs due to FP operations, then any error will initially be introduced in the lower bits. Suppose that both $\mathrm{trainer}$ and $\mathrm{auditor}$ train at a higher FP (e.g., $b_{tr}=64$) precision than the client’s target model precision (e.g., $b_{m}=32$) and then periodically round (e.g., $b_{r}=32$) after intermediate computation steps (e.g., a convolution layer). One might hope that this will “erase" the errors due to nondeterminism, and prevent them from accumulating. Unfortunately, simply rounding to the nearest FP32 after each computation during training is insufficient for determinism. The problem is due to rounding errors that straddle the rounding boundary. Consider Case A in Figure 3, which shows a divergence in the output of a computation using FP64 on two different GPUs. Because the outputs of GPU 1 and 2 are on different sides of the boundary, rounding to the nearest FP32 results in different values, introducing error.
@@
What if the $\mathrm{trainer}$ records their rounding choice (e.g., up, down, none) for every intermediate computation? The $\mathrm{auditor}$ could then copy the $\mathrm{trainer}$’s choice, and therefore round to the exact same value and successfully control for nondeterminism. However, the $\mathrm{auditor}$ should not copy the $\mathrm{trainer}$’s behavior for every output (see Cases B & C, Figure 3). If a computation output on GPU 1 is too close to the rounded value, then it is possible that GPU 2 is also close in distance but from the opposite direction. In this case, the $\mathrm{auditor}$ should ignore the $\mathrm{trainer}$’s choice. We therefore need to introduce a threshold $\tau$ under which the $\mathrm{trainer}$ does not record their rounding choice.
@@
Our method requires upper bounding the divergence $d_{div}$ between any two different GPUs for any intermediate computation $f$ (i.e. difference in outputs for the same input). Let $\epsilon_{b}$ represent the distance between two FP32 values, after rounding to $b_{r}$ bits of the mantissa (Figure 5) and controlling for the exponent. We need to select $b_{r}$ and $\tau$ such that $d_{div}<\epsilon_{b_{r}}$ and $d_{div}<2\tau$ (Figure 3). Because the set of possible FP numbers is finite, there exist optimal bounds for $b_{r}$ and $\tau$. In practice, we find that $b_{r}\leq 32$ and $\tau>0.25\cdot\epsilon_{32}$ are sufficient for standard intermediate computations in neural network training (e.g., convolution, layer norm) in FP64. We study different values for $b_{r}$ in Section 6.
@@ 5.2 Primitives
We assume both $\mathrm{trainer}$ and $\mathrm{auditor}$ train models using the IEEE-754 standard FP numbers (Figure 5). Besides requiring $\mathsf{read}$ and $\mathsf{write}$ disk I/O operations, we define the following functions:
@@
    $\mathsf{rnd}_{b_{r}}(x)$: rounds input $x$ to the nearest FP up to $b_{r}$ bits of the mantissa, as shown in Figure 5.
@@
    $\mathsf{rev}(x,b_{r},c)$: reverses rounding of input $x$ based on logging direction $c$. If $x<\mathsf{rnd}_{b_{r}}(x)$ $c=0$, then return $x$ rounded to the nearest float below $x$ with $b_{r}$ precision. If $x>\mathsf{rnd}_{b_{r}}(x)$ $c=2$, then return $x$ rounded to the nearest float above $x$ with $b_{r}$ precision. Otherwise, do not correct.
@@
    $\mathsf{threshold}(l,b_{r},b_{tr})$: identifies the optimal threshold to log rounding directions (0 or 2) instead of 1, which the $\mathsf{rev}$ function ignores, based on the binary search procedure in Section 5.4.
@@
    $\mathsf{hash}_{\mathsf{sha256}}(\theta)$: creates a SHA-256 hash of provided model weights $\theta$ (in $b_{m}$ precision).
@@
    $\mathsf{tree}(\textit{leaf}_{1},\textit{leaf}_{2}...,\textit{leaf}_{n}):$ create a Merkle tree where each leaf node is the output of $\mathsf{hash}_{\mathsf{sha256}}(\theta)$ for model weights $\theta$ at a given checkpoint, with a checkpointing interval $k$ (Merkle, 1988).
5.3 Training and Auditing
The $\mathrm{trainer}$’s task begins when a client approaches them with dataset $D$, training specifications (epochs $E$, loss function $\mathsf{loss}$, etc.), and a requested model precision $b_{m}$.The $\mathrm{trainer}$ can then choose a training precision $b_{tr}>b_{m}$, a rounding amount $b_{r}\leq b_{m}$, and a checkpointing interval $k$ to periodically store small $\mathsf{hash}_{\mathsf{sha256}}(\theta)$ of model weights $\theta$ in a Merkle tree, for efficient comparison with an eventual $\mathrm{auditor}$. Then, as detailed in Algorithm 1, the $\mathrm{trainer}$ can perform training as normal, but after every intermediate computation (e.g., convolution) perform the $\mathsf{rnd}_{b_{r}}$ operation on each output. Rounding is applied to computations in both the forward and backward passes. Finally, either using a fixed threshold $\tau$ or a layer-specific optimal $\tau$ from the $\mathsf{threshold}$ function described in Section 5.4, the $\mathrm{trainer}$ applies $\mathsf{log}$, which logs rounding choices only for the computations an auditor should copy. The output of the algorithm includes a rounding log file $F$ and the root of the Merkle tree which, along with the shared randomness R and all training parameters, the $\mathrm{trainer}$ can share with any trusted third-pary $\mathrm{auditor}$ who challenges them.
After a client approaches them, the $\mathrm{auditor}$ initiates the verification game described in Section 3. To avoid penalty, the $\mathrm{trainer}$ must cooperate by sharing the rounding amount $b_{r}$, randomness R used in training (e.g., a pseudo-random number generator), the checkpointing interval $k$, and set of rounding logs $F$. The $\mathrm{auditor}$ then follows the training procedure and corrects their rounding choice (e.g., up or down) to match those logged in $F$ using the $\mathsf{rev}$ operation, as detailed in Algorithm 2 (Appendix). By correcting each rounding mismatch during the course of training, the $\mathrm{auditor}$ is able to prevent nondeterminism errors from accumulating. Therefore, the $\mathrm{auditor}$ can store the $\mathsf{hash}_{\mathsf{sha256}}(\theta)$ of model weights $\theta$ in a Merkle tree at interval $k$, knowing that if training was done correctly, the model weights should be identical to the $\mathrm{trainer}$’s at any timestep. The output of Algorithm 2 is the root of the $\mathrm{auditor}$’s Merkle tree, which they can use to compare with the $\mathrm{trainer}$’s root.
5.4 Reducing storage cost
Logging rounding decisions for every neural network layer output during training incurs a large baseline storage cost, and is our main limitation. For dataset $D$, batch size $B$, training epochs $E$, and model layers $L_{\theta}$, the upper bound on the total storage cost for verifiable training with our method is: